RC Meeting and Runplan, Monday, August 27

From HallCWiki
Jump to: navigation, search
  • Status
    • Optics work Sunday did not completely converge - revisit today
    • Initial, rough SHMS steering test Sunday SWING/Monday OWL
    • IHA3H07A - looks like encoder readback issue (cable). Access Tuesday to fix
  • Run plan
    • Monday SWING, Tuesday OWL and DAY
      • Beam optics tuning
      • Big BPM calibration/checkout test plan
      • After initial setup - SHMS steering plan will be executed (please call RC if not already on site). When MCC asks, ramp up magnets (one at a time) to settings for -8.5 GeV.
      • Beam checkout
      • For the following, see detailed run plan here: [1]
        • Detector and trigger checkout. Call experts before start (after hole-target is centered): Malace, Pooser, Berdnikov, Ambrose for detector checkout, Sawatzky, Malace, Pooser, Huber, Trotta for trigger checkout
        • Calibration runs
    • RF recovery scheduled for Tuesday - Hall C going to Controlled Access until ~ 1 pm
    • Controlled Access tasks
      • Harp repair
      • Walk -throughs
      • SHMS PMT investigation